# Configurations, Tessellations and Tone Networks

### v4

@@ VIII Tone Networks and Tessellations
In much of the modern literature of the tonnetz, considerable emphasis has been placed on the role of the 3pp-hexacycles (Cohn’s hexatonic cycles) in understanding voice leading in the analysis of music. Although the significance of this system is undeniable in this respect, one wonders if the attention given to these hexacycles has been at the expense of bypassing the consideration of other structures that are clearly there in the mathematics and hence ought to play a similar role in the associated soundscape. We have given the examples of the $2p$-hexacycles, the $1p$-octacycles, the $3p$-octacycles and the $2p$-decacycles, whose role alongside the four $3p$-hexacycles might be of a comparable level of significance. But there may be other related mathematical chord system structures that are of interest as well.
@@
To emphasize the point, we give a further example of a tone network that can be constructed by purely combinatorial arguments and yet admits a geometric representation that is not unpleasing. Here we consider again the sets of major and minor chords and we construct another map between these sets, this time of a rather different character.
It seems natural that given a major chord one might ask for all the minor chords that can be constructed from it by holding one note fixed and altering the other two. The spirit of this construction is similar to that of the original tonnetz. Thus (as a brief exercise in musical combinatorics will verify), given $CM$ we can construct three minor chords differing from it at two of the tones – namely, $Fm$, GmGm and $C_{\sharp}m$. Similarly, if we are given $Cm$, then we can construct three major chords differing from it at two of the tones – namely, $FM$, GMGM and $BM$.
Figure 17: The Archimedean tonnetze. Each major chord can be changed into a minor chord by shifting two of the tones, and each minor chord can be changed into a major chord by shifting two of the tones. The resulting Levi graph has two connected components, the $CM$ Archimedean tonnetz (on the left) and the $Cm$ Archimedean tonnetz (on the right).
@@
Interestingly, under this new tonnetz the major and minor chords split into two distinct classes. More precisely, the associated bipartite graph is not connected, but rather is given by the union of two distinct connected components. Thus we obtain two sets of twelve chords, each containing six major chords and six minor chords, as shown in Figure 18. Since the presence of any one chord in one of these Archimedean tonnetze determines the rest, we can call the first one the $C$-major Archimedean tonnetz and the other one the $C$-minor Archimedean tonnetz.
@@
As with the other tonnetze that we have considered, one can read off the resulting cycles quite directly from the bipartite graph. There are hexacycles, octacycles and dodecacycles, but there are also tetracycles in this case. The significance of the tetracycles is that these graphs are not the graphs of configurations – for we recall that in order for a bipartite graph to be the Levi graph of a configuration it must be of girth at least six, which has indeed been the case in the previous examples that we have considered. Nonetheless, if we lay out identical copies of the dodecagons in a lattice and connect adjacent vertices across the lattice (rather than across the dodecagon), what emerges is a tessellation of the plane, composed of dodecagons, hexagons, and squares, as in Figure 18.
This tessellation is one of a number known to Kepler (1619). Each dodecagon is surrounded by six hexagons and six squares in a remarkable alternating pattern known as an Archimedean or semi-regular tessellation  [^19] The internal angle of two adjacent sides of a regular $n$-gon is $(n-2)\pi/n$. It follows that if we wish to fit three polygons all with sides of the same length around the same vertex, then the numbers of sides $n_{1},n_{2},n_{3}$ of the polygons must be so that $(n_{1}-2)\pi/n_{1}+(n_{2}-2)\pi/n_{2}+(n_{3}-2)\pi/n_{3}=2\pi$. This condition is satisfied by $n_{1}=4$, $n_{2}=6$, $n_{3}=12$. More generally, there are 21 different solutions with various numbers of polygons and various numbers of sides to fit a collection of regular polygons around a vertex. Of these, 11 can be extended across the plane, giving the 11 so-called Archimedean tilings. of type $\{4,6,12\}$. See, for example, Grünbaum & Shephard (1977, 2016), Conway, Burgiel & Goodman-Strauss (2008) and Wilson (2016). Since the chords at adjacent vertices share exactly one tone in common, we can associate each edge with a tone. The three edges meeting a vertex then give the tones of which the chord corresponding to that vertex is composed. The Hamiltonian cycles in the Archimedean tonnetze ascend in the circle of fifths going counterclockwise around the rim. As a result, two of the three chords a given chord is adjacent to are a fourth or fifth apart.  [^20] A dual tessellation called the ‘Laves tiling’ associated with $\{4,6,12\}$ can be constructed (Grünbaum & Shephard 2016, at pp. 95-98) with the property that the vertices of the original tiling correspond to the tiles of the dual tiling and the tiles of the original tiling correspond to vertices of the dual tiling. The edges of the original tiling and its dual can then be identified. The Laves tilings associated with the Archimedean tonnetze can be seen in Rietsch (2024) at pp. 336-338. Each dodecagon of the original tonnetz corresponds to a vertex of the corresponding Laves tessellation met by twelve edges. Thus, our geometric/combinatorial method gives a construction of the duals of Rietsch’s $G_{2}$ tonnetze, which were obtained with Lie theory.
@@
The Archimedean tonnetze can be used as a basis for composition, and in this respect they are rather useful. In particular, the fact that two notes of a triad change at each transition ensures that the chord sequences are in some respects less ‘bland’ (or trance-like) than those of the cycles of the Eulerian tonnetz, and hence are more interesting to listen to even in the absence of the larger and typically rich tonal environment from which historical examples of the use of the tonnetz are usually extracted. In Figure 19 we illustrate this principle with a short composition in the form of a fanfare based a cycle of the $CM$ Archimedean tonnetz. The presence of the occasional common practice modalities (such as the plagal cadence from $Fm$ to $CM$) embedded in this tonnetz help to moderate the more unusual chord transitions and ground the piece with a deceptive sense of familiar tonality.
@@ IX Dialogo sopra le reti tonali
SALVIATI: – Bolando and Ugopoli. Yes, I can tell you a bit more about it if you wish.
@@
SALVIATI – Well, it’s based on the tonnetz. As you say, they take a mathematical approach. Their central argument, if I interpret it correctly, is that mathematics is not merely a kind of tool box that a music theorist might use for whatever purpose he finds it helpful. The latter would be in the tradition of old-fashioned applied mathematics, but that is not what Bolando and Ugopoli are doing. They treat music and mathematics as equal partners. They look, as I understand it, for certain common structures – mainly of a combinatoric or geometric nature – that arise naturally in both music and mathematics. So yes, to be sure, they are using mathematics to explain certain phenomena in music – but they are also looking for ways in which mathematical structures manifest themselves in music – or how the same underlying idea can work its way out both in music and in mathematics.
SAGREDO – This reminds me slightly of the way that physicists use geometry. The essence of the matter is not simply that geometry can be used as a mathematical tool for analyzing physical phenomena, though it certainly can be used like that, but rather that the physical phenomena are fundamentally geometric in their very nature. Thus, for example, it is not merely the case that space-time can be modelled as a four-dimension manifold with an indefinite metric – rather that space-time really is such a manifold.
SALVIATI – Yes, it’s just like that. So, likewise, they argue, in our considerations of the theory of music, we need to regard the music and the mathematics as a kind of unity. Bolando and Ugopoli begin with Waller’s representation of the tonnetz as a bipartite graph. Waller did not know about the tonnetz, and he had no interest in voice leading and all that, as far as I am aware. He approached the problem of the relationship between major and minor chords from a purely combinatorial and group theoretical point of view.
SAGREDO – Yes, that is what I understand. And he produced that diagram of the relations between major and minor chords in the form of a regular bipartite graph. And now I recall what you were saying earlier – that Bolando and Ugopoli recognized (and seem to have been the first to do so) that this Waller construction is a Levi graph – the graph of a configuration.
@@
SAGREDO – I get it. What Bolando and Ugopoli are suggesting, in essence, is that for every configuration there exists a kind of abstract music – that the points and lines of the configuration correspond to different types of chords (analogous to major and minor), that the adjacency of a point and line correspond to what the music theorists call an efficient voice-leading between the two chords.
SALVIATI – Correct. Bolando and Ugopoli look at a number of examples of such ‘abstract musics’. First they look again at the tonnetz, the usual one. They show that its Levi graph is isomorphic to a certain configuration of twelve lines and twelve points (what we call a $\{12_{3}\}$) with the property that three lines go through each point and three points lie on each line. Such $\{12_{3}\}$ configurations have been investigated since the nineteenth century and it is known that there are 229 of them. Bolando and Ugopoli were able to identify the tonnetz as being the $D222$. This is a configuration that has been studied a lot, along with all 228 of the other $\{12_{3}\}$s, over the last few decades by geometers and combinatorialists, so it is remarkable that no one had previously realized that this is Euler’s tonnetz.
@@
SALVIATI – That is true. And the idea was developed further, later, by the visionary Hugo Riemann, and others, but it was not until the 1980s and 1990s that the theory of the tonnetz took its modern form.
SAGREDO – Well, so I gather. But it is always in the form of those triangular patterns that one sees in Cohn’s book and elsewhere. I can see why the musicologists like them, because they can trace progressions from triangle to triangle, and understand certain aspects of harmony that were previous obscure, which is in fact extremely interesting – but it seems a bit vacuous mathematically speaking, if I may say so, though maybe I am missing something.
@@
SAGREDO – Or even better, one can visualize progressions as movements from point to point in Euclidean space along the straight lines that are joining them in the configuration.
@@
SALVIATI – Not really – even this seemingly narrow definition embraces a vast wealth of material. For other purposes, one looks at configurations in projective spaces such as $\mathbb{RP}^{2}$ or $\mathbb{CP}^{2}$ or higher dimensional projective spaces. There are also certain special projective spaces composed of a finite number of points. These also admit configurations – some of which may be musically interesting. In fact, one can even drop the requirement that the configuration is representable over a field (finite or otherwise) and look at purely combinatorial configurations. The point is that all of these configurations have Levi graphs associated with them on a bijective basis, and so each defines a species of abstract music.
SAGREDO – This is making me dizzy – I think I need another drink. Where is Simplicio? Are you saying that each and every one of these configurations gives rise to a form of music?
@@
SAGREDO – So we can make a list of configurations, and work our through them.
SALVIATI – Yes, let’s do it. First, there is the $\{7_{3}\}$ of Fano, the Levi graph of which has 14 vertices, 7 white and 7 black. It does not admit a representation over the reals or the complexes, but does admit a representation over the finite field $F(2)$.
@@
SALVIATI – Yes. Now, where was I? Hmmm – next we have the Möbius $\{8_{3}\}$, which admits a representation in $\mathbb{CP}^{2}$. And then we have the three $\{9_{3}\}$s, each of which can be realized in $\mathbb{R}^{2}$. The Pappus configuration is among these.
SAGREDO – What about the $\{10_{3}\}$s?
@@
SAGREDO – That does not mean that existing pentatonic music necessarily conforms to the structure of this tonnetz, merely that one can compose new pieces of music by use of it?
@@
SALVIATI – Now, Sagredo, where were we in our list? Well, coming back to the $\{10_{3}\}$s, there is a curious detail that I meant to mention. All but one can be realized over the reals and even over the rationals. But one of them admits no realization. Nonetheless, it is perfectly well-defined as a combinatorial configuration and it has a Levi graph. There is nothing mysterious about it.
@@
SALVIATI – Exactly. So, one has to beware of jumping to conclusions about mathematical objects. Bolando and Ugopoli never claimed that realizability is essential for a configuration to be musically viable. Quite the contrary. The point, rather, is that with that one peculiar exception, all the $\{9_{3}\}$s, $\{10_{3}\}$s, $\{11_{3}\}$s and $\{12_{3}\}$s are realizable.
SAGREDO – So, all but one of the tonnetze that can be constructed for scales of twelve notes or less are realizable over some field, and all but three are realizable over the reals?
SALVIATI – Precisely. So, most of the ones that you will tend to look at in practice, at least in the first instance, will fall into the realizable category. It is not as though one is favouring them in some way. They just happen to be in a heavy majority. But you can make music with any of them, whether or not they are realizable.
@@
SALVIATI – Well, you don’t, in advance. But I think that Coxeter’s remark – interesting configurations are represented by interesting graphs – may be relevant here. The point is that a tonnetz should be regarded as a representation of a configuration. Thus, we may presume, interesting configurations are represented by interesting tonnetze. This is the motivation for studying configurations as a basis for musical analysis. Interesting configurations give rise to interesting musical structures.
@@
SALVIATI – Precisely. That is what our two Academician friends are saying.
SAGREDO – Hmmm, interesting. And what about so-called voice leading. The musicologists make lot of fuss over it, but you have scarcely mentioned it.
SALVIATI – Precisely. Well, Bolando and Ugopoli say that voice leading is a concept that is difficult to define in the abstract. And besides, music gets a bit tedious if the parts only move around in semitones or whole tones, or if chord transitions are all like that.
@@
SALVIATI – Well, they argue that the structure of the tonnetz is already implicit in the relationship between major chords and the notes that make up these chords. No mention of the minor chords is required for this. To be sure, one eventually recovers the standard tonnetz relations between major and minor chords, but these can be derived. One need not make any a priori assumptions about voice leading relations.
SAGREDO That is intellectually satisfying, I suppose. I never bought into all that voice leading dogma.
SALVIATI – Yes, well, you still have your voice leading, but it is implicit in the structure of the incidence relations of the points and lines in the configuration that defines the genre of music in which your piece is composed.
@@
Figure 20: A Tristan genus tonnetz. The twelve minor sixth chords and twelve dominant seventh chords can be arranged in a Levi graph with the incidence geometry of a configuration $\{12_{3}\}$. Each minor sixth is connected to three dominant sevenths and each dominant seventh is connected to three minor sixths. At each transformation two notes are retained in the new chord, one being the dissonant note of the original tetrad, which may or may not be a dissonance in the new one. Edges that cross the tonnetz preserve the dissonant note of the tetrad as a dissonance in the new tetrad. For example, the dissonant $F$ of $G_{\sharp}m^{6}$ is mapped to the dissonant $F$ of $G^{7}$.

### v5

@@ VIII Tone Networks and Tessellations
In much of the modern literature of the Eulerian tonnetz, emphasis has been placed on the role of the 3pp-hexacycles (Cohn’s hexatonic cycles) in understanding voice leading in the analysis of music. Although the significance of this system is undeniable, one wonders if the attention given to these hexacycles has been at the expense of bypassing the consideration of other structures that are there in the mathematics and hence ought to play a role in the associated soundscape. We have given the examples of the $2p$-hexacycles, the $1p$-octacycles, the $3p$-octacycles and the $2p$-decacycles, whose position alongside the four $3p$-hexacycles might be of comparable significance. But there may be other mathematical chord system structures that are of interest as well.
@@
To illustrate the point, we give a further example of a tone network that can be constructed by purely combinatorial arguments and yet admits a geometric representation that is not unpleasing. Here we consider again the sets of major and minor chords and we construct another map between these sets, this time of a different character.
@@
It seems natural that given a major chord one might ask for all the minor chords that can be constructed from it by holding one note fixed and altering the other two. The spirit of this construction is similar to that of the original tonnetz. Thus (as a brief exercise in musical combinatorics will verify), given $CM$ we can construct three minor chords differing from it at two of the tones – namely, $Fm$, GmGm and $C_{\sharp}m$. Similarly, if we are given $Cm$, then we can construct three major chords differing from it at two of the tones – namely, $FM$, GMGM and $BM$.  [^23] The single common tone preserving relations between triads that we have described are known as the ‘obverse’ relations of the usual $P$, $L$ and $R$. For example, $CM\to Fm$ is $L^{\prime}$, the obverse of $L$. These terms were coined by Morris (1998).  Interestingly, under this new tonnetz the major and minor chords split into two distinct classes. More precisely, the associated bipartite graph is not connected, but rather is given by the union of two distinct connected components. Thus we obtain two sets of twelve chords, each containing six major chords and six minor chords, as shown in Figure 19. Since the presence of any one chord in one of these Archimedean tonnetze determines the rest, we can call the first one the $C$-major Archimedean tonnetz and the other one the $C$-minor Archimedean tonnetz.
Figure 19: The Archimedean tonnetze. Each major chord can be changed into a minor chord by shifting two of the tones, and each minor chord can be changed into a major chord by shifting two of the tones. The resulting bipartite graph has two connected components, the $CM$ Archimedean tonnetz (on the left) and the $Cm$ Archimedean tonnetz (on the right).
As with the other tonnetze we have considered, one can read off the resulting cycles directly from the bipartite graph (see Appendix, Table IX). There are hexacycles, octacycles, decacycles and dodecacycles, but also tetracycles such as $\langle Gm$, $F_{\sharp}M$, $C_{\sharp}m$, $CM$, $Gm\rangle$. The significance of the tetracycles is that these graphs are not the graphs of configurations – for in order for a bipartite graph to be the Levi graph of a configuration it must be of girth at least six, which has indeed been the case in the previous examples that we have considered. Nonetheless, if we lay out identical copies of the dodecagons in a lattice and connect adjacent vertices across the lattice (rather than across the dodecagon), what emerges is a tessellation of the plane, composed of dodecagons, hexagons, and squares, as in Figure 20.
@@
This tessellation is one of a number known to Kepler (1619). Each dodecagon is surrounded by six hexagons and six squares in a remarkable alternating pattern known as an Archimedean or semi-regular tessellation  [^24] The internal angle of two adjacent sides of a regular $n$-gon is $(n-2)\pi/n$. It follows that if we wish to fit three polygons all with sides of the same length around the same vertex, then the numbers of sides $n_{1},n_{2},n_{3}$ of the polygons must be so that $(n_{1}-2)\pi/n_{1}+(n_{2}-2)\pi/n_{2}+(n_{3}-2)\pi/n_{3}=2\pi$. This condition is satisfied by $n_{1}=4$, $n_{2}=6$, $n_{3}=12$. More generally, there are 21 different solutions with various numbers of polygons and various numbers of sides to fit a collection of regular polygons around a vertex. Of these, 11 can be extended across the plane, giving the 11 Archimedean tilings. of type $\{4,6,12\}$. See, for example, Grünbaum & Shephard (1977, 2016), Conway, Burgiel & Goodman-Strauss (2008) and Wilson (2016).
Since the chords at adjacent vertices share one tone in common, we can associate each edge with a tone. The three edges meeting a vertex give the tones of which the chord corresponding to that vertex is composed. The Hamiltonian cycles in the Archimedean tonnetze ascend in the circle of fifths going counterclockwise around the rim. As a result, two of the three chords a given chord is adjacent to are a fourth or fifth apart.  [^25] A dual tessellation called the ‘Laves tiling’ associated with $\{4,6,12\}$ can be constructed (Grünbaum & Shephard 2016, at pp. 95-98) with the property that the vertices of the original tiling correspond to the tiles of the dual tiling and the tiles of the original tiling correspond to vertices of the dual tiling. The edges of the original tiling and its dual can then be identified. The Laves tilings associated with the Archimedean tonnetze can be seen in Rietsch (2024) at pp. 336-338. Each dodecagon of the original tonnetz corresponds to a vertex of the corresponding Laves tessellation met by twelve edges. Thus, our geometric/combinatorial method gives a construction of the duals of Rietsch’s $G_{2}$ tonnetze, which were obtained with Lie theory.
@@
The Archimedean tonnetze can be used as a basis for composition, and they are rather useful.
The fact that two notes of a triad change at each transition ensures that the chord sequences are less bland or trance-like than the cycles of the Eulerian tonnetz, and hence are more interesting to listen to even in the absence of the rich tonal environment from which historical examples of the use of the Eulerian tonnetz are usually extracted.  [^26] According to Piston (1985), ‘The purpose of this procedure [the practice of voice leading in harmonic progression] is to insure the smoothest possible connection of two chords, so that one seems to flow into the next. Continued practice of the process will, however, result in rather dull music.’   In Figure 21 we illustrate this point with a short composition in the form of a fanfare based a cycle of the $CM$ Archimedean tonnetz. The occasional common practice modalities embedded in this tonnetz, like the plagal cadence from $Fm$ to $CM$, help to moderate the effects of the more unusual chord transitions and ground the piece with an aura of familiar tonality.  [^27] It is interesting to note how different cycles of the Archimeden tonnetz have contrasting voice-leading properties, along the lines of Tomoczko (2020), pp. 116-128. A counterclockwise movement through the tetracycle $\langle CM,C_{\sharp}m,F_{\sharp}M,Gm,CM\rangle$ induces a one-step scalar transition, whereas movement through a hexacycle returns each voice to its starting pitch, and a clockwise trajectory through the 0pp-dodecacycle leads every voice up an octave. Thus, the figure reflects aspects of the topology of voice-leading spaces. In this respect the tessellation in Figure 20 has an advantage over the associated bipartite graph of Figure 19 since the various cycles are visually more distinct.
@@ IX Dialogo sopra le reti tonali
SALVIATI – Bolando and Ugopoli. Yes, I can tell you a bit more about it if you wish.
@@
SALVIATI – Well, it’s based on the tonnetz. As you say, they take a mathematical approach. Their central argument, if I interpret it correctly, is that mathematics is not merely a kind of tool box that a music theorist might use for whatever purpose he finds it helpful. The latter would be in the tradition of old-fashioned applied mathematics, but that is not what Bolando and Ugopoli are up to. They treat music and mathematics as equal partners. They look for certain common structures – of a combinatoric, geometric or number theoretic nature – that arise naturally in both music and mathematics. So yes, they are using mathematics to explain certain phenomena in music – but they are also looking for ways in which mathematical structures manifest themselves in music – or how the same underlying idea can work its way out both in music and in mathematics.
SAGREDO – This reminds me of the way physicists use geometry. It is not simply that geometry can be used as a tool for analyzing physical phenomena, though it can be used like that, but rather that the physical phenomena are fundamentally geometric in their nature. For example, it is not merely the case that space-time can be modelled as a four-dimensional manifold with an indefinite metric – rather that space-time really is such a manifold.
SALVIATI – Yes, it’s just like that. So, likewise, they argue, in our considerations of the theory of music, we need to regard the music and the mathematics as a kind of unity. Bolando and Ugopoli begin with Waller’s representation of the tonnetz as a bipartite graph. Waller did not know about the tonnetz, and he had no particular interest in voice leading and all that, as far as I am aware. He approached the problem of the relationship between major and minor chords from a purely combinatorial and group theoretical point of view.
SAGREDO – Yes, that is what I understand. He produced that diagram of the relations between major and minor chords in the form of a regular bipartite graph. And I recall what you were saying earlier – that Bolando and Ugopoli recognized (and seem to have been the first to do so) that the Waller construction is a Levi graph – the graph of a configuration.
@@
SAGREDO – I get it. What Bolando and Ugopoli are suggesting is that for every configuration there exists an abstract music – that the points and lines of the configuration represent different types of chords, analogous to major and minor, that the adjacency of a point and line corresponds to what music theorists call efficient voice leading between the chords.
SALVIATI – Correct. Bolando and Ugopoli look at various examples of abstract music systems. First they look at the usual tonnetz. They show that its Levi graph is isomorphic to a configuration of twelve lines and twelve points, what we call a $\{12_{3}\}$, with the property that three lines go through each point and three points lie on each line. Such configurations have been investigated since the 19th century and it is known that there are 229 of them. Bolando and Ugopoli were able to identify the tonnetz as the D222. This configuration has been studied a lot over the last few decades by mathematicians, so it is remarkable that no one had previously realized that it is Euler’s tonnetz.
@@
SALVIATI – True, and the idea was developed further by the visionary Hugo Riemann, but it was not until the 1980s and 1990s that the theory of the tonnetz took its modern form.
SAGREDO – Well, so I gather. But it is always in the form of those triangular patterns that one sees in Cohn’s book and elsewhere. I can see why the musicologists like them, because they can trace progressions from triangle to triangle, and understand certain aspects of harmony that were previous obscure, which is in fact quite interesting – but it seems a bit vacuous mathematically speaking, if I may say so, though maybe I’m missing something.
@@
SAGREDO – Or even better, one can visualize progressions as movements from point to point in Euclidean space along the straight lines that are joining them.
@@
SALVIATI – Not really – even this seemingly narrow definition embraces a wealth of material. For other purposes, one looks at configurations in projective spaces such as $\mathbb{RP}^{2}$ or $\mathbb{CP}^{2}$. There are also certain special projective spaces composed of a finite number of points and these too admit configurations – some of which may be musically interesting. In fact, one can drop the requirement that the configuration is representable over a field and look at purely combinatorial configurations. All of these configurations have Levi graphs associated with them and each defines a species of abstract music.
SAGREDO – This is making me dizzy – I think I need another drink. Where is Simplicio? Are you saying to me that each and every one of these configurations gives rise to a form of music?
@@
SAGREDO – So we can make a list of configurations, and work our way through them.
SALVIATI – Yes, let’s do it. First, there is the $\{7_{3}\}$ of Fano, the Levi graph of which has 14 vertices, 7 white and 7 black. It does not admit a representation over the reals or the complexes, but does admit a geometric representation over the finite field $\mathbb{F}(2)$.
@@
SALVIATI – Correct. Now, where was I? Hmmm – next we have the Möbius $\{8_{3}\}$, which admits a geometric representation in $\mathbb{CP}^{2}$. And then we have the three $\{9_{3}\}$s, each of which can be realized in $\mathbb{R}^{2}$. The Pappus configuration is among these.
SAGREDO – What about the $\{10_{3}\}$ configurations?
@@
SAGREDO – That does not mean that existing pentatonic music always conforms to the structure of this tonnetz, merely that one can compose new pieces of music by use of it?
@@
SALVIATI – Now, Sagredo, where were we in our list? Well, coming back to the $\{10_{3}\}$ configurations, there is one very curious detail that I meant to mention. All but one can be realized over the reals and even over the rationals. But one of them admits no such realization. Nonetheless, it is perfectly well-defined as a combinatorial configuration and it has a Levi graph. There is nothing mysterious about it.
@@
SALVIATI – Exactly. So, one has to beware of jumping to conclusions about mathematical objects and their properties simply on the basis of the names given to them. Bolando and Ugopoli never claimed that realizability is essential for a configuration to be musically viable. Quite the contrary. The point, rather, is that with that one peculiar exception, all the $\{7_{3}\}$, $\{8_{3}\}$, $\{9_{3}\}$, $\{10_{3}\}$, $\{11_{3}\}$ and $\{12_{3}\}$ configurations are realizable.
SAGREDO – So, all but one of the 270 tonnetze that can be constructed for scales of twelve notes or less are geometrically realizable over some field, and all but three of them are realizable in two-dimensional Euclidean space?
SALVIATI – Precisely. Nearly all of the ones that you will tend to look at in practice, at least in the first instance, fall into the ‘geometrically realizable’ category. It is not as though one is favouring or cherry-picking them in some way. They just happen to be in a heavy majority. But you can make music with any of these 270 tonnetze, whether or not they are geometrically realizable.
@@
SALVIATI – Well, you don’t, in advance. But I think that Coxeter’s remark – ‘interesting configurations are represented by interesting graphs’ – may be relevant here. The point is that a tonnetz should be regarded as a representation of a configuration. It follows, therefore, we may presume, that interesting configurations are represented by interesting tonnetze. This is the motivation for studying configurations as a basis for musical analysis. Interesting configurations give rise to interesting musical structures.
@@
SALVIATI – Precisely. That is exactly what our two Academician friends are saying: interesting configurations give rise to interesting musical structures – in particular, to interesting tonnetze.
SAGREDO – Hmmm, fascinating. And what about voice leading. The musicologists make lot of fuss over it, but you have scarcely mentioned it.
SALVIATI – Precisely. Well, Bolando and Ugopoli say that voice leading is a concept that is difficult to define in the abstract. In any case, music gets a bit tedious if the parts only move around in semitones or whole tones, or if the chord transitions are all like that.
@@
SALVIATI – They argue that the structure of the tonnetz is implicit in the relation between major chords and the notes that make up the chords. No mention of the minor chords is needed. One eventually recovers the usual tonnetz relations between major and minor chords, but these are derived. No a priori assumptions about voice leading are required.
SAGREDO – That is intellectually satisfying, I suppose. I never bought into all that voice leading dogma. Isn’t it really a secondary matter, at least mathematically speaking?
SALVIATI – Yes, well, you can say that. You still have your voice leading in the Bolando-Ugopoli system, but it is implicit in the structure of the incidence relations of the points and lines in the configuration that defines the genre of music in which your piece is composed.